OpenStudy (anonymous):

Please place the following in y=mx+b format. No decimals please also place all fractions as improper and in lowest terms.Also please explan how you got the answer. y-7=1/4(x-3)

OpenStudy (anonymous):

y-7=1/4x - 3/4 y=1/4x+27/4

OpenStudy (anonymous):

1) distribute the 1/4 to x and -3 2) add 7 to both sides

OpenStudy (anonymous):

thnx but isnt it y=1/4+27/4

OpenStudy (anonymous):

but thnx cus u helped me get the concept

OpenStudy (anonymous):

* sorry y=1/4+25/4